OpenStudy (anonymous):

express the relationship between a small change in x and the corresponding change in y in the form dy=f'(x)dx f(x)=2-acosx

OpenStudy (amistre64):

is acos short for arccos? looks like a programing shorthand

OpenStudy (amistre64):

if so then let y = acos(x) cos(y) = x -sin(y) y' = 1 y' = -csc(y) , and we know y = acos(x) y' = -csc(acos(x))